# Improving variational counterdiabatic driving with weighted actions and computer algebra

### v1

@@ Introduction
for $0\leq\lambda\leq 1$, where $\Lambda_{\mathrm{NN}}$ is the set of the nearest-neighbor pairs of spins ($(i,j)$ and $(j,i)$ are counted only once). We set $\hbar=1$ in all numerical results. We use the two-dimensional square lattice of width $N_{\mathrm{w}}$ and height $N_{\mathrm{h}}$ with open boundary condition. The number of spins is $N=N_{\mathrm{w}}N_{\mathrm{h}}$, and the number of nearest-neighbor pairs is $|\Lambda_{\mathrm{NN}}|=N_{\mathrm{w}}(N_{\mathrm{h}}-1)+(N_{\mathrm{w}}-1)N_{\mathrm{h}}$. The transverse magnetic field has a fixed strength $1.0$. The longitudinal magnetic field $h_{i}$ is sampled from an i.i.d. gamma distribution with mean $1.0$ and standard deviation $0.5$, which ensures $h_{i}\geq 0$. By changing the sign and coefficients of the coupling term, we realize ferromagnetic, antiferromagnetic, and spin-glass systems. For the ferromagnetic system, we choose the minus sign from the $\pm$ sign and sample $J_{ij}$ from an i.i.d. gamma distribution with mean $1.0$ and standard deviation $0.5$. For the antiferromagnetic system, we choose the plus sign and draw $J_{ij}$ from the same gamma distribution. For the spin-glass system, we sample $J_{ij}$ from an i.i.d. Gaussian distribution with mean 00 and standard deviation $1.0$ (the choice of the $\pm$ sign is irrelevant).
@@
Figure 1: Determination of the driving coefficients using the weighted variational method with different values of $K$, exemplified with a ferromagnetic Ising model with $N=12$. (a) The additional weight $w_{n}^{(K)}(\lambda)$ on the eigenstates plotted against the energy eigenvalues $\epsilon_{n}(\lambda)$ for $K=1,\dots,5$ (purple–yellow). We use $\lambda=0.25$ as an example. Each weight is normalized so that the area under the curve is constant for better visualization, as the overall normalization is unimportant. The arrows show the energy shift $E_{\lambda}^{(K)}$ for $K=2,\dots,5$, which corresponds to the center of the weight function. The energies $\epsilon_{1}$ and $\epsilon_{D}$ are the lowest and the highest energy eigenvalues, and the gray histogram shows the density of states. (b) The driving coefficients $\alpha_{\mu}(\lambda)$ for $\mu=1,\dots,12$ of the one-body driving obtained from the weighted variational method with $K=1,\dots,5$.
We determine the driving coefficients $\bm{\alpha}^{(K)}(\lambda)$ by the weighted variational method in Eq. (21) with degrees $K=1,\dots,5$. We then numerically solve the Schrödinger equation in Eq. (1) with the Hamiltonian $H(\lambda_{t})+\dot{\lambda_{t}}V(\bm{\alpha}^{(K)}(\lambda_{t}))$ and the initial state $|\phi_{1}(0)\rangle$ using the QuTiP package on Python [75, 76]. We write the resulting time evolution as $|\psi^{(K)}(t)\rangle$. For comparison, we also compute the time evolution without the driving term by setting $V(\lambda_{t})=0$ in Eq. (1), whose result is written as $|\psi^{(\varnothing)}(t)\rangle$. We use a simple linear schedule, $\lambda_{t}=t/t_{\mathrm{d}}$ for $0\leq t\leq t_{\mathrm{d}}$.
@@
The gain measures the relative increase of the fidelity compared to the conventional method, and it trivially satisfies $\mathcal{G}_{\mathrm{f}}^{(1)}=1$. These performance measures depend on the realization of the random parameters $\{h_{i}\}$ and $\{J_{ij}\}$, the choice of the driving terms (one-body or two-body), the system size $N$, and the protocol duration $t_{\mathrm{d}}$ in addition to the degree $K$, but we make these dependencies implicit for brevity.
Our algorithm for computing the driving coefficients $\bm{\alpha}^{(K)}(\lambda)$ is efficient, and it can be applied to systems with a large number of spins. For example, ten minutes calculation can deal with $N=3400$ for $K=2$, $N=360$ for $K=3$, $N=84$ for $K=4$, and $N=37$ for $K=5$ with our C++ implementation available at Ref. [69]. These values are obtained from the calculations of the driving coefficients for a one-dimensional Ising model ($N_{\mathrm{h}}=1$ and $N_{\mathrm{w}}=N$ in the above setup) with the one-body driving at 100 points of $\lambda\in[0,1]$ using the algorithm for the entire range of $\lambda$. We used an ordinary laptop computer (MacBook Pro with Apple M3 Pro chip and 18 GB memory) and the Clang compiler with the level-O2 compiler optimization and CPU parallelization.
Despite this efficiency, we will focus entirely on smaller systems, $N_{\mathrm{h}}=3$ and $N_{\mathrm{w}}=3,4,5$, and hence $N=9,12,15$ in the following demonstration. This choice is to avoid the exponentially increasing cost of the numerical simulation of larger quantum systems.
@@
We first demonstrate how the weighted variational method with different degrees results in different driving coefficients using a ferromagnetic system with $N=12$ as an example. Figure 1(a) shows the weight associated with the $n$th eigenstate, $w_{n}^{(K)}(\lambda)$, for $K=1,\dots,5$. The weight is more concentrated on low-energy states as $K$ increases. In Fig. 1(b), we plot the resulting driving coefficients $\alpha_{\mu}^{(K)}(\lambda)$ for $K=1,\dots,5$ for the one-body driving ansatz. As seen from the figure, both the shape and the magnitude of the driving coefficients vary with the degree $K$. A higher degree method tends to result in a larger magnitude in this specific system, while this tendency is not always true for other systems. The increase in the maximum magnitude is at most three times for almost all cases. Thus, the difficulty of experimental implementation does not qualitatively increase.
@@
We show the performance of the weighted variational method in the same ferromagnetic system of $N=12$ with the one-body driving in Fig. 2. Figure 2(a) shows the time evolution of the fidelity $\mathcal{F}^{(K)}(t)$ with a short duration $t_{\mathrm{d}}=0.01$. The fidelity drops from unity to almost zero ($\simeq 2.4\times 10^{-4}$) in the absence of the driving term ($K=\varnothing$). The final fidelity remains around $0.089$ with the driving protocol calculated by the conventional method ($K=1$). With our weighted variational method ($K\geq 2$), the final fidelity is significantly improved up to around $0.80$ at $K=5$, which is $8.9$ times larger than the conventional method.
The final fidelity $\mathcal{F}_{\mathrm{f}}^{(K)}$ depends on the protocol duration $t_{\mathrm{d}}$, as shown in Fig. 2(b). When the duration is sufficiently long, the time evolution with and without CD driving both achieve fidelities close to unity. In this regime, $\dot{\lambda}_{t}=1/t_{\mathrm{d}}$ is small, and the CD driving term becomes negligible in the Schrödinger equation in Eq. (3). Therefore, the unit fidelity is understood from the adiabatic theorem [70]. As the duration $t_{\mathrm{d}}$ becomes shorter, the fidelity decreases and converges to a constant value that depends on $K$. In this regime, $\dot{\lambda}_{t}=1/t_{\mathrm{d}}$ is large, and the first term $\kappa_{n}(t)c_{n}(t)$ becomes negligible in Eq. (3). We can then recast Eq. (3) into a form independent of the duration $t_{\mathrm{d}}$ by using the chain rule $(\dot{\lambda}_{t})^{-1}\partial_{t}c_{n}(t)=\partial_{\lambda}c_{n}(t_{\lambda})$, where $t_{\lambda}$ is the inverse function of $\lambda_{t}$. This makes the final fidelity independent of $t_{\mathrm{d}}$ [27, 67]. In other words, the time evolution in this short-duration limit is determined solely by the discrepancy between the exact and approximated AGPs. In the following, we always focus on this regime and set $t_{\mathrm{d}}=0.01$ to investigate the impact of the approximation error.
@@
In Fig. 3, we show that this enhancement of the final fidelity is not limited to the specific instance (the set of realization of the random parameters $\{h_{i}\}$ and $\{J_{ij}\}$) but holds more generally. We generate 100 instances of ferromagnetic systems for each of the system sizes $N=9,12,15$, calculate the driving protocols of the one-body driving for $K=1,\dots,5$ for each instance, and perform a quantum simulation for every $K\in\{\varnothing,1,\dots,5\}$ for each instance. As shown in Fig. 3(a), the final fidelity $\mathcal{F}_{\mathrm{f}}^{(K)}$ varies from instance to instance, but it is in the same order of magnitude. The median of the final fidelity increases as the degree $K$ increases for every system size $N$. In Fig. 3(b), we show the gain $\mathcal{G}_{\mathrm{f}}^{(K)}$, i.e., the relative increase of the final fidelity. The gain is greater than unity in all instances and for all $K\geq 2$, and the median gain increases as we increase $K$ for every system size $N$. For $K=5$, the median of the gain $\mathcal{G}_{\mathrm{f}}^{(5)}$ is about $4.9$, $8.7$, and $14$ for $N=9,12,15$, respectively, demonstrating a significant improvement in the final fidelity.
@@
Figure 3 also shows the system-size dependence of the results, where the final fidelity and gain follow different trends. The median final fidelity $\mathcal{F}_{\mathrm{f}}^{(K)}$ decreases for a larger system size $N$ for each $K\in\{\varnothing,1,\dots,5\}$. This behavior in the absence of CD driving ($K=\varnothing$) and with the conventional method ($K=1$) is well known in the literature (see Ref. [29] for a detailed investigation), and it reflects the intrinsic—and presumably unavoidable—difficulty of realizing adiabatic processes with a limited number of driving fields. This fundamental difficulty persists in the weighted variational method, as suggested by the results with $K\geq 2$. On the other hand, the median gain $\mathcal{G}_{\mathrm{f}}^{(K)}$ increases as the size $N$ increases for every fixed $K\geq 2$. This result implies that the weighted variational method becomes more advantageous for larger systems in terms of the gain. It suggests that our proposal is still effective for even larger systems than those tested here.
@@
Figure 4: Numerical test of the weighted variational method in antiferromagnetic systems. We generate $100$ instances for every system size $N=9,12,15$ and perform simulations similarly to Fig. 3. (a) Final fidelity $\mathcal{F}_{\mathrm{f}}^{(K)}$ with the one-body driving, varying the system size $N$ and the degree $K$. (b) Final fidelity $\mathcal{F}_{\mathrm{f}}^{(K)}$ with the two-body driving. (c) Gain $\mathcal{G}_{\mathrm{f}}^{(K)}$ with the two-body driving. In all panels, some outliers are above the upper bound of the axes.
@@
Figure 5: Numerical test of the weighted variational method in spin-glass systems. We generate $100$ instances for every system size $N=9,12,15$ and apply the weighted variational method with the one-body driving. (a) Final fidelity $\mathcal{F}_{\mathrm{f}}^{(K)}$ with varying the system size $N$ and the degree $K$. (b) Gain $\mathcal{G}_{\mathrm{f}}^{(K)}$. In both panels, some outliers are above the upper bound of the axes.
@@
The ferromagnetic system presented in the previous section has no frustration or competition between the coupling and external fields. To test the general applicability of the weighted variational method to more disordered and chaotic systems, we perform a similar numerical test on antiferromagnetic and spin-glass systems. We randomly generate 300 instances of antiferromagnetic systems and 300 instances of spin-glass systems (100 each for $N=9,12,15$), compute the driving coefficients, and perform quantum simulations similarly to the previous section.
@@
In the antiferromagnetic systems with the one-body driving, conventional CD driving does harm to the adiabatic process, but the weighted variational method can assist the process [Fig. 4(a)]. The final fidelity with the conventional method $\mathcal{F}_{\mathrm{f}}^{(1)}$ is smaller than the fidelity without CD driving $\mathcal{F}_{\mathrm{f}}^{(\varnothing)}$ in most instances (287 instances out of 300). Nevertheless, the median of $\mathcal{F}_{\mathrm{f}}^{(K)}$ increases as $K$ increases for all system sizes, and the final fidelity $\mathcal{F}_{\mathrm{f}}^{(5)}$ at $K=5$ is greater than $\mathcal{F}_{\mathrm{f}}^{(\varnothing)}$ for about 75% of the instances. Thus, the weighted variational method can recover the advantage of CD driving even when the conventional variational method does not lead to improvement.
@@
We can circumvent this peculiar behavior of antiferromagnetic systems by using the two-body driving [Figs. 4(b) and (c)]. As shown in Fig. 4(b), the final fidelity with the conventional method $\mathcal{F}_{\mathrm{f}}^{(1)}$ is larger than the final fidelity without CD driving $\mathcal{F}_{\mathrm{f}}^{(\varnothing)}$ in most instances (283 instances out of 300), which implies that the conventional method is advantageous when used with the two-body driving. The final fidelity is further improved by the weighted variational method with $K\geq 2$. As shown in Fig. 4(c), the gain $\mathcal{G}_{\mathrm{f}}^{(K)}$ exceeds unity in all instances for all $K\geq 2$. The median gain reaches $5.3,9.9,11$ for $N=9,12,15$, respectively, at $K=5$.
@@
For the spin-glass systems, the one-body driving is valid for a large fraction of the instances, while the instance-to-instance variations of $\mathcal{F}_{\mathrm{f}}^{(K)}$ and $\mathcal{G}_{\mathrm{f}}^{(K)}$ are stronger than for the previous two classes of systems (Fig. 5). The conventional method $\mathcal{F}_{\mathrm{f}}^{(1)}$ is better than the bare time evolution $\mathcal{F}_{\mathrm{f}}^{(\varnothing)}$ for a large fraction of instances (259 instances out of 300), and the weighted variational method with a larger degree $K$ further improves the final fidelity. The gain is greater than unity for all $K\geq 2$ in most of the instances (295 instances out of 300). Overall, the gain is more dramatic than that in the ferromagnetic and antiferromagnetic systems. The median of $\mathcal{G}_{\mathrm{f}}^{(5)}$ is about $10,13,19$ for $N=9,12,15$, respectively.
@@
In all cases above, a larger system size results in a larger median gain $\mathcal{G}_{\mathrm{f}}^{(K)}$ for every fixed $K\geq 2$ [Figs. 4(c) and 5(b)]. This fact suggests that our method is still powerful for chaotic and disordered systems of larger sizes.
Analyses of the driving coefficients
@@
We analyze the properties of the driving coefficients obtained from the weighted variational method and discuss how these properties contribute to the observed improvement in the final fidelity, thereby bridging between the theoretical framework in Secs. II and III and the numerical results in Sec. IV. The weighted action extends the conventional action in two aspects, as discussed in Sec. II.2: it assigns more weights to important matrix elements and incorporates nonlocal terms. We examine how these features manifest in the resulting driving coefficients from both theoretical and numerical analyses. Throughout this section, we fix a single $\lambda$ and omit the $\lambda$-dependence of quantities such as $H(\lambda)$, $\epsilon_{n}(\lambda)$, $\mathcal{S}_{\lambda}^{(K)}[V]$, $w_{mn}^{(K)}(\lambda)$, $w_{n}^{(K)}(\lambda)$, $\bm{\alpha}^{(K)}(\lambda)$, and $E_{\lambda}^{(K)}$ except for Eq. (34).
@@
We analyze how the weights on low-energy states in weighted actions change the driving coefficients and, hence, improve the fidelity. We decompose $\mathcal{S}^{(K)}[V]$ into contributions from each eigenstate $\mathcal{T}^{(K,n)}[V]$, which we call the partial action (see Appendix A.5 for derivation):

### v3

@@ Introduction
for $0\leq\lambda\leq 1$, where $\Lambda_{\mathrm{NN}}$ is the set of the nearest-neighbor pairs of spins ($(i,j)$ and $(j,i)$ are counted only once). We set $\hbar=1$ in all numerical results. We use the two-dimensional square lattice of width $N_{\mathrm{w}}$ and height $N_{\mathrm{h}}$ with open boundary condition. The number of spins is $N=N_{\mathrm{w}}N_{\mathrm{h}}$, and the number of nearest-neighbor pairs is $|\Lambda_{\mathrm{NN}}|=N_{\mathrm{w}}(N_{\mathrm{h}}-1)+(N_{\mathrm{w}}-1)N_{\mathrm{h}}$. The longitudinal magnetic field $h_{i}$ is sampled from an i.i.d. gamma distribution with mean $1.0$ and standard deviation $0.5$, which ensures $h_{i}\geq 0$. By changing the sign and coefficients of the coupling term, we realize ferromagnetic, antiferromagnetic, and spin-glass systems. For the ferromagnetic system, we choose the minus sign from the $\pm$ sign and sample $J_{ij}$ from an i.i.d. gamma distribution with mean $1.0$ and standard deviation $0.5$. For the antiferromagnetic system, we choose the plus sign and draw $J_{ij}$ from the same gamma distribution. For the spin-glass system, we sample $J_{ij}$ from an i.i.d. Gaussian distribution with mean 0 and standard deviation $1.0$ (the choice of the $\pm$ sign is irrelevant).
@@
Figure 2: Determination of the driving coefficients using the weighted variational method with different values of $K$, exemplified with a ferromagnetic Ising model with $N=12$. (a) The weight on the eigenstates $w_{n}^{(K)}(\lambda)$ plotted against the energy eigenvalues $\epsilon_{n}(\lambda)$ for $K=1,\dots,5$ (purple–yellow). We use $\lambda=0.25$ as an example. Each weight is normalized so that the area under the curve is constant for better visualization, as the overall normalization is unimportant. The arrows show the energy shift $E_{\lambda}^{(K)}$ for $K=2,\dots,5$, which corresponds to the center of the weight function. The energies $\epsilon_{1}$ and $\epsilon_{D}$ are the lowest and the highest energy eigenvalues, and the gray histogram shows the density of states. (b) The driving coefficients $\alpha^{(K)}_{\mu}(\lambda)$ at two representative sites of the one-body driving obtained from the weighted variational method with $K=1,\dots,5$. Solid curves represent $\alpha^{(K)}_{\mu}(\lambda)$ of one of the sites, and dashed curves are for another site. We plot the coefficients for all sites in Fig. S1 in Supplemental Material [82].
We determine the driving coefficients $\bm{\alpha}^{(K)}(\lambda)$ by the weighted variational method in Eq. (21) with degrees $K=1,\dots,5$, and we numerically simulate the time evolution. We take a simple linear schedule, $\lambda_{t}=t/t_{\mathrm{d}}$ for $0\leq t\leq t_{\mathrm{d}}$, where $t_{\mathrm{d}}$ is the protocol duration, and solve the Schrödinger equation in Eq. (1) with Hamiltonian $H(\lambda_{t})+\dot{\lambda_{t}}V(\bm{\alpha}^{(K)}(\lambda_{t}))$ and the initial state $|\phi_{1}(0)\rangle$ using the QuTiP package in Python [83, 84]. We write the resulting time evolution as $|\psi^{(K)}(t)\rangle$. For comparison, we also compute the time evolution without the driving term by setting $V(\lambda_{t})=0$ in Eq. (1), whose result is written as $|\psi^{(\varnothing)}(t)\rangle$.
@@
The gain measures the relative increase of the fidelity compared to the conventional method, and it trivially satisfies $\mathcal{G}_{\mathrm{f}}^{(1)}=1$. These performance measures depend not only on degree $K$ but on the realization of the random parameters $\{h_{i}\}$ and $\{J_{ij}\}$, the choice of driving terms (one-body or two-body), the system size $N$, and the protocol duration $t_{\mathrm{d}}$, but we make these dependencies implicit for brevity.
Our algorithm for computing the driving coefficients $\bm{\alpha}^{(K)}(\lambda)$ is efficient, and it can be applied to systems with a large number of spins. For example, ten minutes of calculation can deal with $N=4330$ for $K=2$, $N=360$ for $K=3$, $N=84$ for $K=4$, and $N=37$ for $K=5$ with our C++ implementation available in Ref. [73]. These values are obtained from the calculations of the driving coefficients for a one-dimensional Ising model ($N_{\mathrm{h}}=1$ and $N_{\mathrm{w}}=N$ in the above setup) with the one-body driving at 100 values of $\lambda\in[0,1]$. We used Algorithm 4 in Appendix C.4 for the algebraic calculation and the conjugate gradient method to solve Eq. (18). We used an ordinary laptop computer (MacBook Pro with Apple M3 Pro chip and 18 GB memory) and the Clang compiler with the level-O2 compiler optimization and CPU parallelization.
Despite this efficiency, we will focus entirely on smaller systems, $N_{\mathrm{h}}=3$ and $N_{\mathrm{w}}=3,4,5$, and hence $N=9,12,15$ in the following demonstration. The system size is constrained by the cost of the numerical simulation for our demonstrative purposes, which would not be necessary when testing our method on real quantum systems.
@@
We first demonstrate how the weighted variational method with different degrees results in different driving coefficients using a ferromagnetic system with $N=12$ as an example. Figure 2(a) shows the weight associated with the $n$th eigenstate, $w_{n}^{(K)}(\lambda)$, for $K=1,\dots,5$. The weight is more concentrated on low-energy states as $K$ increases. In Fig. 2(b), we plot the resulting driving coefficients $\alpha_{\mu}^{(K)}(\lambda)$ for $K=1,\dots,5$ for the one-body driving ansatz. As seen from the figure, both the shape and the magnitude of the driving coefficients vary with the degree $K$. A higher degree method tends to result in a larger magnitude in this specific system, while this tendency is not always true for other systems (see Fig. S1 in Supplemental Material [82]). The increase in the maximum magnitude is at most three times for almost all cases. Thus, the difficulty of experimental implementation does not qualitatively increase.
@@
We test the weighted variational method in the same ferromagnetic system of $N=12$ with the one-body driving, which results in significantly improved fidelity. Figure 3(a) shows the time evolution of the fidelity $\mathcal{F}^{(K)}(t)$ with a short duration $t_{\mathrm{d}}=0.01$. The fidelity drops from unity to almost zero ($\simeq 2.4\times 10^{-4}$) in the absence of the driving term ($K=\varnothing$). The final fidelity remains around $0.089$ with the driving protocol calculated by the conventional method ($K=1$). With our weighted variational method ($K\geq 2$), the final fidelity is significantly improved up to around $0.80$ at $K=5$, which is $8.9$ times larger than the conventional method.
The final fidelity $\mathcal{F}_{\mathrm{f}}^{(K)}$ depends on the protocol duration $t_{\mathrm{d}}$, as shown in Fig. 3(b). When the duration is sufficiently long, the time evolution with and without CD driving both achieve fidelities close to unity. In this regime, $\dot{\lambda}_{t}=1/t_{\mathrm{d}}$ is small, and the CD driving term becomes negligible in the Schrödinger equation in Eq. (3). Then, the unit fidelity is understood from the adiabatic theorem [74]. As the duration $t_{\mathrm{d}}$ becomes shorter, the fidelity decreases and converges to a constant value that depends on $K$. In this regime, $\dot{\lambda}_{t}=1/t_{\mathrm{d}}$ is large, and the first term $\kappa_{n}(t)c_{n}(t)$ becomes negligible in Eq. (3). We can then recast Eq. (3) into a form independent of the duration $t_{\mathrm{d}}$ by using the chain rule $(\dot{\lambda}_{t})^{-1}\partial_{t}c_{n}(t)=\partial_{\lambda}c_{n}(t_{\lambda})$, where $t_{\lambda}$ is the inverse function of $\lambda_{t}$. This makes the final fidelity independent of $t_{\mathrm{d}}$ [27, 71]. In other words, the time evolution in this short-duration limit is determined solely by the discrepancy between the exact and approximated AGPs. In the following, we always focus on this regime and set $t_{\mathrm{d}}=0.01$ to investigate the impact of the discrepancy.
@@
Figure 4 shows that the enhancement of the final fidelity is not limited to the specific instance (the set of realization of the random parameters $\{h_{i}\}$ and $\{J_{ij}\}$) but holds more generally. We generate 100 instances of ferromagnetic systems for each of the system sizes $N=9,12,15$, calculate the driving protocols of the one-body driving for $K=1,\dots,5$ for each instance, and perform a quantum simulation for every $K\in\{\varnothing,1,\dots,5\}$ for each instance. As shown in Fig. 4(a), the final fidelity $\mathcal{F}_{\mathrm{f}}^{(K)}$ varies from instance to instance, but it is within the same order of magnitude. The median of the final fidelity increases as the degree $K$ increases for every system size $N$. In Fig. 4(b), we show the gain $\mathcal{G}_{\mathrm{f}}^{(K)}$, i.e., the relative increase of the final fidelity. The gain is greater than unity in all instances and for all $K\geq 2$, and the median gain increases as we increase $K$ for every system size $N$. For $K=5$, the median of the gain $\mathcal{G}_{\mathrm{f}}^{(5)}$ is about $4.9$, $8.7$, and $14$ for $N=9,12,15$, respectively, demonstrating a significant improvement in the final fidelity.
@@
Figure 4 also shows the system-size dependence of the results, where the final fidelity $\mathcal{F}_{\mathrm{f}}^{(K)}$ and gain $\mathcal{G}_{\mathrm{f}}^{(K)}$ follow different trends. The median final fidelity decreases for a larger system size $N$ for each $K\in\{\varnothing,1,\dots,5\}$ [Fig. 4(a)]. This behavior in the absence of CD driving ($K=\varnothing$) and with the conventional method ($K=1$) is well known in the literature (see Ref. [29] for a detailed investigation), and it reflects the intrinsic—and presumably unavoidable—difficulty of realizing adiabatic processes with a limited number of driving fields. This fundamental difficulty persists in the weighted variational method, as suggested by our results with $K\geq 2$. On the other hand, the median gain increases as the size $N$ increases for every fixed $K\geq 2$ [Fig. 4(b)]. This result implies that the weighted variational method becomes more advantageous for larger systems in terms of gain. It suggests that our proposal is still effective for even larger systems than those tested here.
@@
Figure 5: Numerical test of the weighted variational method in antiferromagnetic systems. We generate $100$ instances for every system size $N=9,12,15$ and perform simulations similarly to Fig. 4. (a) Final fidelity $\mathcal{F}_{\mathrm{f}}^{(K)}$ with the one-body driving, varying the system size NN and the degree $K$. (b) Final fidelity $\mathcal{F}_{\mathrm{f}}^{(K)}$ with the two-body driving. (c) Gain $\mathcal{G}_{\mathrm{f}}^{(K)}$ with the two-body driving. In all panels, some outliers are above the upper bound of the axes.
@@
Figure 6: Numerical test of the weighted variational method in spin-glass systems. We generate $100$ instances for every system size $N=9,12,15$ and apply the weighted variational method with the one-body driving similarly to Fig. 4. (a) Final fidelity $\mathcal{F}_{\mathrm{f}}^{(K)}$ with varying the system size NN and the degree $K$. (b) Gain $\mathcal{G}_{\mathrm{f}}^{(K)}$. In both panels, some outliers are above the upper bound of the axes.
@@
The ferromagnetic system presented in the previous section has no frustration or competition between the coupling and external fields. To test the general applicability of the weighted variational method to more disordered and chaotic systems, we perform a similar numerical test on antiferromagnetic and spin-glass systems. We randomly generate 300 instances of antiferromagnetic systems and 300 instances of spin-glass systems (100 each for $N=9,12,15$), compute the driving coefficients, and perform quantum simulations similarly to the previous section. We plot representative driving coefficients in Fig. S1 in Supplemental Material [82], which tends to show more complicated behavior than in Fig. 2(b).
We first observe the time-dependent fidelity $\mathcal{F}^{(K)}(t)$ and the dependence of the final fidelity on protocol duration $t_{\mathrm{d}}$ for representative systems (Fig. S2 in Supplemental Material [82]), confirming that they are qualitatively similar to those in Fig. 3. Concretely, the fidelity $\mathcal{F}^{(K)}(t)$ drops from unity to a lower value over time evolution, and when we shorten the duration $t_{\mathrm{d}}$, the final fidelity $\mathcal{F}_{\mathrm{f}}^{(K)}$ converges to a constant value dependent on $K$. Thus, we can evaluate the overall performance by focusing on the final fidelity with the shortest duration $t_{\mathrm{d}}=0.01$, as we have done for ferromagnetic systems.
@@
In the antiferromagnetic systems with the one-body driving, conventional CD driving does harm to the adiabatic process, but the weighted variational method can assist the process [Fig. 5(a)]. The final fidelity with the conventional method $\mathcal{F}_{\mathrm{f}}^{(1)}$ is smaller than the fidelity without CD driving $\mathcal{F}_{\mathrm{f}}^{(\varnothing)}$ in most instances (287 instances out of 300), meaning that the conventional method hinders the adiabatic process. Nevertheless, the median of $\mathcal{F}_{\mathrm{f}}^{(K)}$ increases as $K$ increases for all system sizes, and the final fidelity $\mathcal{F}_{\mathrm{f}}^{(5)}$ at $K=5$ is greater than $\mathcal{F}_{\mathrm{f}}^{(\varnothing)}$ for about 75% of the instances. Thus, the weighted variational method can recover the advantage of CD driving even when the conventional variational method does not lead to improvement.
@@
We can circumvent this peculiar behavior of antiferromagnetic systems by using the two-body driving [Figs. 5(b) and (c)]. The final fidelity with the conventional method $\mathcal{F}_{\mathrm{f}}^{(1)}$ is higher than the final fidelity without CD driving $\mathcal{F}_{\mathrm{f}}^{(\varnothing)}$ in most instances (283 instances out of 300), which implies that the conventional method is advantageous when used with the two-body driving. The final fidelity is further improved by the weighted variational method with $K\geq 2$ [Fig. 5(b)]. As shown in Fig. 5(c), the gain $\mathcal{G}_{\mathrm{f}}^{(K)}$ exceeds unity in all instances for all $K\geq 2$. The median gain reaches $5.3,9.9,11$ for $N=9,12,15$, respectively, at $K=5$.
@@
For the spin-glass systems, the one-body driving is valid for a large fraction of instances, while the instance-to-instance variations of $\mathcal{F}_{\mathrm{f}}^{(K)}$ and $\mathcal{G}_{\mathrm{f}}^{(K)}$ are stronger than for the previous two classes of systems (Fig. 6). The conventional method $\mathcal{F}_{\mathrm{f}}^{(1)}$ is better than the bare time evolution $\mathcal{F}_{\mathrm{f}}^{(\varnothing)}$ for a large fraction of instances (259 instances out of 300), and the weighted variational method with a larger degree $K$ further improves the final fidelity [Fig. 6(a)]. The gain is greater than unity for all $K\geq 2$ in most of the instances (295 instances out of 300), as shown in Fig. 6(c). The gain is more dramatic than that in the ferromagnetic and antiferromagnetic systems. The median of $\mathcal{G}_{\mathrm{f}}^{(5)}$ is about $10,13,19$ for $N=9,12,15$, respectively.
@@
In all the cases above, a larger system size results in a larger median gain $\mathcal{G}_{\mathrm{f}}^{(K)}$ for every fixed $K\geq 2$ [Fig. 5(c) and 6(b)]. This fact suggests that our method remains powerful for chaotic and disordered systems of larger sizes, even beyond the sizes tested in this section.
Properties of the driving coefficients from the weighted variational method
@@
We analyze the properties of the driving coefficients obtained from the weighted variational method and discuss how these properties contribute to the observed improvement in the final fidelity, thereby bridging between the theoretical framework in Secs. II and III and the numerical results in Sec. IV. As discussed in Sec. II.3, the weighted action extends the conventional action in two aspects: it assigns more weights to important matrix elements and incorporates nonlocal terms. We examine how these features manifest in the resulting driving coefficients from both theoretical and numerical analyses. Throughout this section, we fix a single $\lambda$ and omit the $\lambda$-dependence of quantities such as $H(\lambda)$, $\epsilon_{n}(\lambda)$, $\Phi(\lambda)$, $\mathcal{S}_{\lambda}^{(K)}[V]$, $w_{mn}^{(K)}(\lambda)$, $w_{n}^{(K)}(\lambda)$, $\bm{\alpha}^{(K)}(\lambda)$, and $E_{\lambda}^{(K)}$, except in Eq. (33).
@@
We analyze how the weights on low-energy states affect the action and improve the ground-state fidelity. We invoke the expansion of the weighted action into matrix elements [Eq. (14)]. We will show that the matrix elements associated with lower-energy states are closely related to the ground-state fidelity by using the quantum speed limit formula and an additional numerical calculation. This theoretical analysis supports the expectation that emphasizing these elements is favorable for improving the ground-state fidelity.
We decompose $\mathcal{S}^{(K)}[V]$ into contributions from each eigenstate $\mathcal{T}^{(K,n)}[V]$, which we call the partial action (see Appendix B.4 for derivation):